OpenStudy (anonymous):

The perimeter of a rectangle is fifty centimeters. The length of the rectangle is twenty centimeters. What is the width of the rectangle?

hartnn (hartnn):

Perimeter is just twice the sum of length and width P =2 (L+W) here, P = 50, L=20 just find W

OpenStudy (nurali):

W=x L=x+9 P=L+W+L+W 50cm=(x+9cm)+(x)+(x+9cm)+(x) 50cm=4x+18cm 50cm-18cm=4x 32cm=4x 32cm/4=4x/4 8cm=x So we know W=x, and L=x+9, we also know that the width is 8 cm so the length is 8+9, which is 17 cm

OpenStudy (amistre64):

not real sure how Nuralis setup relates. But it seems to be the solution to a different question from: http://answers.yahoo.com/question/index?qid=20110307200029AAlZqae